### v1

# SHAM-OT: Rapid Subhalo Abundance Matching with Optimal Transport

Abstract
Subhalo abundance matching (SHAM) is widely used for connecting galaxies to dark matter haloes. In SHAM, galaxies and (sub-)haloes are sorted according to their mass (or mass proxy) and matched by their rank order. In this work, we show that SHAM is the solution of the optimal transport (OT) problem on empirical distributions (samples or catalogues) for any metric transport cost function. In the limit of large number of samples, it converges to the solution of the OT problem between continuous distributions. We propose SHAM-OT: a formulation of abundance matching where the halo-galaxy relation is obtained as the optimal transport plan between galaxy and halo mass functions. By working directly on these (discretized) functions, SHAM-OT eliminates the need for sampling or sorting and is solved using efficient OT algorithms at negligible compute and memory cost. Scatter in the galaxy-halo relation can be naturally incorporated through regularization of the transport plan. SHAM-OT can easily be generalized to multiple marginal distributions. We validate our method using analytical tests with varying cosmology and luminosity function parameters, and on simulated halo catalogues. The efficiency of SHAM-OT makes it particularly advantageous for Bayesian inference that requires marginalization over stellar mass or luminosity function uncertainties.
keywords:
cosmology: large-scale structure of Universe – galaxies: luminosity function, mass function – methods: numerical
pubyear: 2025pagerange: SHAM-OT: Rapid Subhalo Abundance Matching with Optimal TransportSHAM-OT: Rapid Subhalo Abundance Matching with Optimal Transport
1 Introduction
Dark matter N-body simulations have become an essential tool for understanding the large scale distribution of matter in our universe and its evolution. However, since these simulations only model dark matter, linking their results to observable galaxies requires the understanding of the galaxy-halo connection (see Wechsler & Tinker (2018) and references therein). A commonly used method to populate the haloes from these simulations with galaxies is subhalo abundance matching (SHAM, Kravtsov et al., 2004; Conroy et al., 2006; Vale & Ostriker, 2004). SHAM provides a simple framework for establishing the galaxy-halo connection, while maintaining the necessary complexity through its treatment of substructure. The SHAM solution is obtained by sorting the galaxy and halo catalogues by their mass (or mass proxy, such as the absolute magnitude) and matching them by their rank order. Its success in accurately describing this relationship has been demonstrated through multiple applications to observational data (Conroy et al., 2006; Behroozi et al., 2010; Guo et al., 2010; Moster et al., 2010; Wetzel & White, 2010; Trujillo-Gomez et al., 2011; Watson et al., 2012; Nuza et al., 2013; Reddick et al., 2013; Simha et al., 2012; Chaves-Montero et al., 2016; Berner et al., 2024).
To create the SHAM matching function, the galaxy and halo distributions are represented by samples; a large number of samples is required for the matching to be accurate. This procedure is typically repeated at multiple redshifts. The generation and sorting of these samples may become computationally expensive for large catalogues, in terms of both time and memory. While this is manageable for single simulations, it may become a bottleneck for Bayesian inference that marginalizes over uncertainties in the stellar mass or luminosity functions, as the matching has to be re-computed for any new combination of parameters. Another important element of SHAM is the scatter in the galaxy-halo relation; its inclusion should preserve the halo mass and galaxy luminosity functions. For certain models, such as log-normal stellar mass scatter (Behroozi et al. (2010), hereafter B10), the “original” stellar mass function is computed by deconvolving its “noisy” counterpart with the scatter model.
In this work, we present SHAM-OT: a novel framework that reformulates the SHAM problem as optimal transport (OT). Generally, OT finds a 2D joint probability distribution (or a transport plan) such that its marginals are equal to two given 1D distributions, while minimising the cost of transporting one into the other, according to a given cost function (see Villani, 2009, for review). We show that the SHAM procedure is exactly the same as the solution to the optimal transport problem of empirical distributions (represented by samples), under certain assumptions on the transport cost function. Framing SHAM as OT allows us to operate directly in the large sample limit and use continuous PDF representations of the distributions, instead of samples. Thus, standard OT solvers can be used to find the matching from discretised halo and galaxy mass functions, without the need of creating catalogues. In this work we show the benefits of using the SHAM-OT framework: (i) computational efficiency, (ii) natural inclusion of scatter in the matching relation via OT plan regularization, and (iii) straightforward extension to matching more than two distributions using multi-marginal OT.
2 Optimal Transport
We briefly introduce the optimal transport (OT) framework, using the Monge-Kantorovich formulation (Kantorovich, 1942; Villani, 2003); see Villani (2009) for a general overview. Intuitively, OT is a way to transport mass from one distribution to another while minimizing the cost of transport.
Let’s consider continuous optimal transport. For measurable spaces $X$ and $Y$, let $a\in\mathcal{P}(X)$ and $b\in\mathcal{P}(Y)$ be probability measures on $X$ and $Y$, respectively. A transport plan $\pi$ is a joint probability measure on $X\times Y$ whose marginals on $X$ and $Y$ equal $a$ and $b$, respectively. The set of all possible transport plans for $a$ and $b$ is
$\footnotesize\Pi(a,b)\!=\!\Bigl{\{}\,\pi$ $\in\!\mathcal{P}(X\times Y):$ (1)
$\int_{A\times Y}\pi(da\,dy)\!=\!a(A),\ \int_{X\times B}\pi(dx\,db)\!=\!b(B)\Bigr{\}},$
for all $A\!\subseteq\!X,B\!\subseteq\!Y$ that are also measurable. Given a cost function $c\!:\!X\times Y\!\to\!\mathbb{R}_{+}$, the Kantorovich OT problem finds the optimal transport plan $\pi^{*}$ that minimizes the total cost
$\pi^{*}=\arg\min_{\pi\in\Pi(a,b)}\int_{X\times Y}c(x,y)\,\pi(dx\,dy).$ (2)
Now let’s consider the discrete optimal transport. For two discrete distributions $\mathbf{p}\!\in\!\mathbb{R}_{+}^{n}$ and $\mathbf{q}\!\in\!\mathbb{R}_{+}^{m}$, normalized to $\mathbf{p}^{\top}\mathbf{1}_{n}\!=\!\mathbf{q}^{\top}\mathbf{1}_{m}\!=\!1$, where $\mathbf{1}$ is a vector of ones, the transport plan $\mathbf{Q}\!\in\!\Gamma(\mathbf{p},\mathbf{q})$ is a doubly stochastic matrix:
$\Gamma(\mathbf{p},\mathbf{q})=\{\mathbf{Q}\in\mathbb{R}_{+}^{n\times m}:\mathbf{Q}\mathbf{1}_{n}=\mathbf{p},\mathbf{Q}^{\top}\mathbf{1}_{m}=\mathbf{q}\}.$ (3)
[…]
$\mathbf{Q}^{*}=\arg\min_{\mathbf{Q}\in\Gamma(\mathbf{p},\mathbf{q})}\langle\mathbf{C},\mathbf{Q}\rangle.$ (4)
This solution may not be unique. The problem can be expressed as a linear program and solved using standard solvers. The transport plan can be penalized by a regularization term, which encourages smoothness of the transport plan and causes the solution to be unique. The most common penalty is entropic regularization, where a negative entropy term $H(\mathbf{Q})\!=\!\langle\mathbf{Q},\ln\mathbf{Q}\rangle$ is added:
$\mathbf{Q}^{*}_{\epsilon}=\arg\min_{\mathbf{Q}\in\Gamma(\mathbf{p},\mathbf{q})}\langle\mathbf{C},\mathbf{Q}\rangle+\epsilon H(\mathbf{Q}),$ (5)
In the limit of $\epsilon\!\to\!0$, the solution converges to that of the unregularized OT (Cominetti & Martín, 1994). Entropic OT has the additional advantage that it can be solved efficiently using the Sinkhorn algorithm (Sinkhorn, 1964). The Sinkhorn method in Alg. 1 is an iterative algorithm that alternates between row and column normalization. Its simplicity and parallelizability makes OT applicable to large-scale problems (Cuturi, 2013).
Algorithm 1 Sinkhorn Algorithm
0:  Cost matrix $\mathbf{C}$, marginals $\mathbf{p},\mathbf{q}$, regularization $\epsilon$
1:  $\mathbf{K}\leftarrow\exp(-\mathbf{C}/\epsilon)$Compute Gibbs kernel}
2:  $\mathbf{u}\leftarrow\mathbf{1}$Initialize scaling vector}
3:  repeat
4:     $\mathbf{v}\leftarrow\mathbf{q}/(\mathbf{K}\mathbf{u})$Scale columns}
5:     $\mathbf{u}\leftarrow\mathbf{p}/(\mathbf{K}^{\top}\mathbf{v})$Scale rows}
6:  until convergence
7:  return  $\text{diag}(\mathbf{u})\mathbf{K}\text{diag}(\mathbf{v})$Compute transport plan}
Figure 1: Analytical halo mass and luminosity function and the corresponding galaxy-halo connection derived using SHAM-OT and standard SHAM at redshift $z\!=\!0.5$. The standard SHAM approach can only be applied up to a halo mass of approximately $10^{14}M_{\odot}$ due to limited number of massive objects in the sampled catalogue, which here contains $\sim 10^{8}$ objects. In contrast, SHAM-OT operates directly on marginal functions, removing the need for sampling and sorting.
3 The Subhalo Abundance Matching Problem
Figure 2: Analytical galaxy-halo relation and their dependence on different parameters at redshift $z\!=\!0.5$. We show the impact of the matter density $\Omega_{m}$ (top left), the Hubble parameter $h$ (top right), the amplitude of the luminosity function at redshift $z\!=\!0$ given by $\phi_{1}^{*}$ (bottom left) and $M_{2}^{*}$ which is responsible for the redshift evolution of $M^{*}$, therefore changing the redshift evolution of the knee of the luminosity function.
Abundance matching techniques find the relationship between galaxies and dark matter haloes. In their simplest form, these methods assume a monotonic relationship between the halo mass and a given galaxy property, such as stellar mass or absolute magnitude. The galaxy-halo connection is then obtained by ranking both galaxies and haloes by these properties, and matching the most massive (or brightest) galaxies with the most massive haloes. Subhalo abundance matching (SHAM) extends this method to account for subhaloes and produces a relation that agrees better with observations (Kravtsov et al., 2004; Conroy et al., 2006; Vale & Ostriker, 2004). Additionally, more realistic models incorporate scatter in the galaxy-halo relation (Behroozi et al., 2010; Trujillo-Gomez et al., 2011; Reddick et al., 2013; Zentner et al., 2014). For example, B10 demonstrates how log-normal scatter of stellar mass at a fixed halo mass can be added, while preserving the marginal distribution of the stellar mass function. Furthermore, differentiating between passive red galaxies and star-forming blue galaxies enhances modelling accuracy, as these populations exhibit distinct preferences for occupation in haloes versus subhaloes (Dragomir et al., 2018; Rodríguez-Puebla et al., 2015; Yamamoto et al., 2015; Berner et al., 2024).
The core computational challenge is the same for all the SHAM models listed above: sorting two catalogues to establish the galaxy-halo relation for each redshift interval. The accuracy of matching relies on using very large catalogues, possibly with billions of objects, which becomes computationally intensive in terms of computing time and memory footprint, especially if re-computed frequently.
4 SHAM as Optimal Transport
We show that the SHAM procedure can be expressed as optimal transport, and provide a sketch of the proof using basic results from OT and statistics theory. Let $\{x_{i}\}_{N}{\sim}a$ and $\{y_{i}\}_{N}{\sim}b$ be i.i.d. samples from two distributions. In the case of SHAM, these correspond to the halo and the galaxy catalogues, respectively. An empirical distribution $\pi_{N}^{\sigma}\in\mathcal{P}(X\times Y)$ created by pairing the samples is
$\pi_{N}^{\sigma}=\frac{1}{N}\sum_{i=1}^{N}\delta(x_{i},y_{\sigma(i)}),$ (6)
where $\sigma{\in}I!$ is a permutation sequence of the indices $I{=}\{1,\ldots,N\}$, and $\delta$ is a Dirac measure. Let us define the metric cost function $C(x,y)$ with the metric properties: $C(0,0)\!=\!0,C(x,y)\!\geq\!0,C(x,y)\!=\!C(y,x),C(x,y)\!\leq\!C(x,z)\!+\!C(z,y)\ \forall\ x\!\in\!X,y\!\in\!Y,z\!\in\!Z.$ The goal is to find a permutation $\sigma$ that minimizes this cost, similarly to Eqn. 2
$\pi_{N}^{*}=\arg\min_{\sigma\in I!}\int_{X\times Y}C(x,y)\pi_{N}^{\sigma}(dx\,dy).$ (7)
Let $x^{*}_{k},y^{*}_{k}$ be the $k$-th order statistics, such that the samples are sorted in ascending order $x_{1}^{*}\!\leq\!x_{2}^{*}\!\leq\!\cdots\!\leq\!x_{N}^{*},\ y_{1}^{*}\!\leq\!y_{2}^{*}\!\leq\!\cdots\!\leq\!y_{N}^{*},$ For the metric cost $C(\cdot,\cdot)$, the optimal transport plan $\pi_{N}^{*}$ is given by matching the order statistics (Bobkov & Ledoux, 2019, Lemma 4.2):
$\pi_{N}^{*}=\frac{1}{N}\sum_{k=1}^{N}\delta(x_{k}^{*},y_{k}^{*})\in\mathcal{P}(X\times Y),$ (8)
Intuitively, for the sorted matched samples, the cost cannot be lowered by exchanging any two matches. Theorem 4.3 in Bobkov & Ledoux (2019) establishes the approximation bounds for transport cost between the empirical distribution and the true distribution. By the Fundamental Theorem of Statistics (Glivenko–Cantelli), the empirical distribution converges to the true distribution $\pi^{*}_{N}\to\pi^{*}$ when $N\to\infty$ almost surely. Therefore, in the limit of $N\to\infty$, the solution to the SHAM procedure is equal to the solution of the continuous optimal transport in Eqn. 2 with any metric cost function. The regularized OT, however, will depend on the choice of the cost.
Figure 3: Comparison of scatter models: SHAM-OT with entropic regularization and standard SHAM. Left: Mean magnitude as a function of halo mass for different regularization $\epsilon$ for SHAM-OT and scatter parameters $\sigma_{\mathrm{mag}}$ for the standard SHAM. Right: Conditional transport plan for $\epsilon\!=\!3\!\times\!10^{-6}$ and $\sigma_{\mathrm{mag}}\!=\!0.6$. The cost is defined using $\mu\!\propto\!\log_{10}M$ and $\nu\!\propto\!-\!m$ (see Sec. 5).
Figure 4: Impact of different cost functions on the conditional transport plan, for quadratic (left) and absolute difference (right) costs, with $\epsilon\!=\!0.01$. Quadratic cost results in increased scatter for lower masses. Absolute difference cost reduces the scatter for low masses. This demonstrates how regularization can serve as a flexible tool to model different types of noise behaviour.
5 Matching halo mass and galaxy luminosity with Optimal Transport
We demonstrate how optimal transport can be used to derive the galaxy-halo relation from analytical distributions. The halo mass function gives the number of haloes $n_{\mathrm{h}}$ per mass $M$
$\small\frac{\mathrm{d}n_{\mathrm{h}}}{\mathrm{d}M}=f(\sigma)\frac{\bar{\rho}}{M}\frac{\mathrm{d}\ln\sigma^{-1}}{\mathrm{d}M},$ (9)
where $\sigma$ is the variance of a spherical top hat containing the mass $M$, $\bar{\rho}$ is the mean matter density and $f(\sigma)$ is the halo multiplicity function. We use the multiplicity function by Watson et al. (2013) which is given by
$\small f(\sigma)=A\left[\left(\frac{\beta}{\sigma}\right)^{\alpha}+1\right]\exp\left(-\frac{\gamma}{\sigma^{2}}\right),$ (10)
with the parameters $A\!=\!0.282$, $\alpha\!=\!2.163$, $\beta\!=\!1.406$, and $\gamma\!=\!1.210$. The subhalo mass function, for a subhalo of mass $M_{\mathrm{s}}$, is given by
$\frac{\mathrm{d}n}{\mathrm{d}\ln(M_{\mathrm{s}}/M)}=\frac{f_{0}}{\beta\Gamma(1-\gamma)}\left(\frac{M_{\mathrm{s}}}{\beta M}\right)^{-\gamma}\exp\left(-\frac{M_{\mathrm{s}}}{\beta M}\right),$ (11)
where $M$ is the central halo mass, and the parameters are $\beta\!=\!0.3$, $\gamma\!=\!0.9$, and $f_{0}\!=\!0.1$, following Mo et al. (2010). The subhalo mass function is computed as a weighted average of the subhalo mass functions over the entire range of central halo masses considered, using the halo mass function as the weight at each mass. This yields the full subhalo mass function, which is then combined with the halo mass function to derive the dark matter host mass function.
Popular choices for the galaxy mass proxy include the stellar mass and the absolute magnitude (see Wechsler & Tinker (2018) for an overview). In this work, we use the absolute magnitude in the B-band in the parametrization from Fischbacher et al. (2024a). The number of galaxies $n_{\mathrm{g}}$ per absolute magnitude $m$ is given by
$\small\frac{\mathrm{d}n_{\mathrm{g}}}{\mathrm{d}m}=\frac{2}{5}\ln(10)\phi^{*}(z)10^{\frac{2}{5}(M^{*}(z)-M)(\alpha+1)}\exp\left(-10^{\frac{2}{5}\left(M^{*}\left(z\right)-M\right)}\right),$ (12)
where $\alpha$ is a free parameter and $\phi^{*}(z)$ and $M^{*}(z)$ are functions of redshift parametrized by
$\small\phi^{*}(z)=\phi^{*}_{1}\exp(\phi^{*}_{2}z),\ \ M^{*}(z)=\begin{cases}M^{*}_{1}+M^{*}_{2}\log(1+z)&\text{if }z<z_{0}\\ M^{*}_{1}+M^{*}_{2}\log(1+z_{0})&\text{if }z\geq z_{0}\end{cases},$
where $M^{*}_{1},M_{2}^{*},\phi^{*}_{1},\phi^{*}_{2}$ and $z_{0}$ are free parameters. The free parameters are fixed to the first index of the blue population of the publicly available model by Fischbacher et al. (2024a).
The transport of distributions requires them to have the same total mass. We ensure this by defining a lower limit on the host mass: we choose $M^{\mathrm{min}}\!=\!10^{10}\!M_{\odot}$, and then compute the necessary faintest absolute magnitude such that the number of haloes and galaxies per unit volume is equal. For numerical reasons, we also define an upper limit on both the halo mass ($M^{\mathrm{max}}{=}10^{19}M_{\odot}$) and the brightest absolute magnitude ($m^{\mathrm{min}}$=-27). These numbers are chosen such that the number of haloes or galaxies above those limits is so small that it does not impact the matching results from SHAM. We create a linearly-spaced vector of negative magnitudes $\tilde{\bm{\nu}}\!\in\!\mathbb{R}^{n}\!:\!\tilde{\nu}_{i}\!=\!-m_{i}$ and logarithmically-spaced halo mass vector $\tilde{\bm{\mu}}\!\in\!\mathbb{R}^{n}\!:\!\tilde{\mu}_{i}\!=\!\log_{10}M_{i}$, both with $n\!=\!500$ within the limits described above, resulting in sampling spacing of $\Delta\tilde{\nu}\!=\!0.036$ and $\Delta\tilde{\mu}\!=\!0.018$. We generate the discretized halo mass function $\mathbf{p}\!\in\!\mathbb{R}_{+}^{n}$ using equations 9 and 11, and the galaxy absolute magnitude function $\mathbf{q}\in\mathbb{R}_{+}^{n}$ using Eqn. 12. We show $\mathbf{p}$ and $\mathbf{q}$ in side panels in Fig. 1. To calculate the cost, we define $\bm{\nu}\!=\!(\tilde{\bm{\nu}}\!-\!\tilde{\nu}_{\mathrm{ref}})/\Delta\tilde{\nu}$ and $\bm{\mu}\!=\!(\tilde{\bm{\mu}}\!-\!\tilde{\mu}_{\mathrm{ref}})/\Delta\tilde{\mu}$, with reference values $\tilde{\nu}_{\mathrm{ref}}\!=\!9.98$ and $\mu_{\mathrm{ref}}\!=\!9.09$. We define the cost matrix $\mathbf{C}$ as $C_{ij}\!=\!(\nu_{i}\!-\!\mu_{j})^{2}/\mu_{j}^{2}$. We obtain the optimal transport plan by solving the problem in Eqn. 4 using linear programming. In the standard SHAM approach, we sample catalogues in a redshift interval $z\!\in\![0.495,0.505]$, which is typical for practical abundance matching, resulting in $\sim\!10^{8}$ objects. The galaxy-halo connection can only be computed up to a mass limit of about $10^{14}M_{\odot}$, as the rapid decline of the mass functions prevents any higher-mass haloes in the simulated catalogue.
The resulting optimal transport plan $\mathbf{Q}$ is presented in Fig. 1 and compared with the standard SHAM approach: they align almost perfectly. The residual differences stem from the fact that the standard SHAM is affected by noise from finite number of pairs, especially at the high-mass end. SHAM-OT overcomes this limitation by acting directly on the discretized functions, which can be extended to arbitrarily large halo masses. Furthermore, we find a computational speed improvement of $O(100)$ for SHAM-OT compared to standard SHAM. SHAM-OT’s computational efficiency allows us to easily study the impact of changes in the luminosity or halo mass function. In Fig. 2, we show how different values of the cosmological and luminosity function parameters impact the galaxy-halo connection.
Figure 5: Galaxy-halo relation for red and blue galaxies in a redshift range from $z\!=\!0$ to $z\!=\!1.5$. The relatively abrupt change in the relation at $8\!\times\!10^{12}M_{\odot}$ is due to the sharp mass limit used for the red/blue assignment.
6 Entropic Regularization as a Scatter Model
In SHAM, the scatter noise is often modelled by a conditional probability distribution on the stellar mass given a fixed halo mass (B10). A common choice is a log-normal distribution, the scale of which may vary with halo mass. This noise is added after the matching procedure. To ensure the preservation of the stellar mass distribution after adding the noise, the “direct” stellar mass function is used for matching, which is created by deconvolving the “true” stellar mass function with the desired noise model (B10). This deconvolution step may cause difficulties at high-mass end with low number of objects, as well as for more advanced noise models.
In the OT framework, the marginals of the transport plan are guaranteed to match the input distributions. The effect of scatter can be modelled naturally by the use of regularization of the transport plan: increasing the regularization parameter $\epsilon$ will result in a smoother plan. Various noise model behaviours, including the dependence on halo mass or galaxy luminosity, can be designed through different choices of the cost function, as long as it satisfies the metric requirements, as discussed in Sec. 4.
We demonstrate the capacity of the entropic regularization to model various scattering properties, including ones that mimic the log-normal approach. We use the same analytical prescription for the halo mass and luminosity function as described in Sec. 5. We compute the standard SHAM solution with Gaussian scatter in magnitude for three noise levels, using the method from B10 for three values of scatter $\sigma_{\mathrm{mag}}$. In SHAM-OT, we choose the cost matrix $C_{ij}\!=\!(\nu_{i}\!-\!\mu_{j})^{2}/\mu_{j}^{2}$. Three values of $\epsilon$ and the form of the cost function were chosen so that the resulting scatter closely resembles the Gaussian scatter of the standard approach. As before, we obtain the OT solution with Sinkhorn (Alg. 1). Fig. 4 shows the comparison between the two approaches for noise levels. SHAM-OT matches the log-normal model well, both in terms of the mean of the distribution (left panel) and its shape (right panel). It reproduces the trend that higher scatter shifts the mean transport plan from high halo masses to fainter galaxies and finds similar probability distributions.
Beyond emulating the log-normal scatter, regularized OT offers a flexible framework to model different scatter behaviours. Fig. 4 illustrates this flexibility by comparing the conditional transport plan for two different choices of cost functions. Further exploration of cost functions, along with the development of their physically motivation, is left for future work.
7 SHAM-OT with haloes from simulations
In practice, the halo mass function is often not analytically given, but can instead be computed from empirical halo distributions based on simulations. We use the approximate code PINOCCHIO (Monaco et al., 2002a; Taffoni et al., 2002; Monaco et al., 2002b, 2013; Munari et al., 2017; Rizzo et al., 2017) to simulate the past light cone of dark matter haloes. PINOCCHIO uses Lagrangian perturbation theory as an approximation which makes it faster than full N-body simulations. We assign subhaloes to the halo catalogue based on the merger tree using the method described in Berner et al. (2022) and label the haloes and subhaloes as hosts for red or blue galaxies based on the model from Berner et al. (2024). This model is based on two key parameters: the halo mass limit $M_{\text{limit}}\!=\!8\!\times\!10^{12}\,h^{-1}M_{\odot}$ and the satellite quenching time $t_{\text{quench}}\!=\!2.0$ Gyr. Central haloes with masses above $M_{\text{limit}}$ host red galaxies, while lower-mass haloes host blue galaxies. Galaxies in subhaloes remain blue until they have spent more than $t_{\text{quench}}$ within their host, after which they transition to red. As a result, we obtain catalogues of haloes and subhaloes with their corresponding mass, position in 3D space, and population label (i.e. red or blue). Note that the methods above can easily be replaced by any other approach that provides a halo catalogue containing the same parameters. We split the host catalogue into redshift bins with width $\Delta z\!=\!0.01$. For each bin, we obtain a host mass function as a histogram, including the upper (heaviest host in the redshift bin) and lower (mass resolution of the simulation) mass limits. We then utilize the luminosity function as defined in Fischbacher et al. (2024a) with the parameters for the red and blue population from the first index of the posterior distribution. Similar to the procedure described in Sec. 5, we determine the faintest absolute magnitude required such that the number of hosts equals the expected number of galaxies in this redshift bin. Using the constructed distributions for the host mass and galaxy absolute magnitudes, we compute the OT solution for each redshift bin, for both red and blue galaxies. The resulting galaxy-halo relation at different redshifts, for both red and blue galaxies, is shown in Fig. 5.
8 Matching more than two functions with multi-marginal optimal transport
Multi-marginal optimal transport is a generalization of optimal transport that applies to any number of marginal distributions (see e.g. Benamou et al., 2015; Pass, 2015). We will consider three distributions in the discrete setting, but the method can be extended analogously to higher dimensions. Given distributions $\mathbf{p}\!\in\!\mathbb{R}_{+}^{n},\mathbf{q}\!\in\!\mathbb{R}_{+}^{m},\mathbf{r}\!\in\!\mathbb{R}_{+}^{k}$ such that $\mathbf{p}^{\top}\mathbf{1}\!=\mathbf{q}^{\top}\mathbf{1}\!=\mathbf{r}^{\top}\mathbf{1}\!=\!1$ and a cost tensor $\mathbf{C}\!\in\!\mathbb{R}_{+}^{n\times m\times k}$, entropy-regularized multi-marginal OT finds the transport plan $\mathbf{Q}\!\in\!\mathbb{R}_{+}^{n\times m\times k}$ that minimizes the cost
$\min_{\mathbf{Q}\in\Gamma(\mathbf{p},\mathbf{q},\mathbf{r})}\langle\mathbf{C},\mathbf{Q}\rangle+\epsilon H(\mathbf{Q}),$ (13)
where the set of allowed transport plans is
$\Gamma(\mathbf{p},\mathbf{q},\mathbf{r})\!=\!\Bigl{\{}\mathbf{Q}\!\in\!\mathbb{R}_{+}^{n\times m\times k}:\sum_{jk}Q_{ijk}\!=\!p_{i},\sum_{ik}Q_{ijk}\!=\!q_{j},\sum_{ij}Q_{ijk}\!=\!r_{k}\Bigl{\}}$
The Sinkhorn algorithm, extended to this setting (Benamou et al., 2015), is shown in Alg. 2. We demonstrate the multi-marginal matching between three distributions: halo mass, galaxy luminosity, and HI mass. The HI mass function is given by
$\frac{\mathrm{d}n_{\mathrm{HI}}}{\mathrm{d}M}=\ln(10)\phi_{*}\left(\frac{M_{\mathrm{HI}}}{M_{*}}\right)^{\alpha+1}\exp\left(-\frac{M_{\mathrm{HI}}}{M_{*}}\right)$ (14)
with $M_{*}\!=\!10^{9.94},\phi_{*}\!=\!4.5\times 10^{-3}\,\mathrm{Mpc}^{-3}\mathrm{dex}^{-1}$ and $\alpha\!=\!-1.25$ following Jones et al. (2018). Similar to before, we make sure that the lower mass limit of HI mass is chosen such that the three matched distributions are normalized consistently. The multi-marginal matching plan is shown in Fig. 6. The 1D marginals are preserved by construction.
Figure 6: Conditional distributions from the multi-marginal optimal transport plan $Q$ with halo mass $p$, absolute magnitude $q$, and HI mass $r$, using $\epsilon{=}0.001$. Density panels show conditional transport plans $\sum_{k}Q_{ijk}/p_{i}$, $\sum_{i}Q_{ijk}/q_{j}$, $\sum_{j}Q_{ijk}/r_{k}$. The 1D marginals are preserved by the transport plan by construction.
0:  Cost tensor $\mathbf{C}$, marginals: $\mathbf{p}$, $\mathbf{q}$, $\mathbf{r}$, regularization: $\epsilon$
  $\mathbf{K}\leftarrow\exp(-\mathbf{C}/\epsilon)$Compute Gibbs kernel}
  $\mathbf{u}\leftarrow\mathbf{1},\mathbf{v}\leftarrow\mathbf{1},\mathbf{w}\leftarrow\mathbf{1}$Initialize scaling vectors}
  repeat
$u_{i}$ $\leftarrow p_{i}/\sum_{jk}K_{ijk}v_{j}w_{k}$
$v_{j}$ $\leftarrow q_{j}/\sum_{ik}K_{ijk}u_{i}w_{k}$
$w_{k}$ $\leftarrow r_{k}/\sum_{ij}K_{ijk}u_{i}v_{j}$
  until convergence
  return  $Q_{ijk}=K_{ijk}u_{i}v_{j}w_{k}$Compute transport plan}
Algorithm 2 Sinkhorn algorithm for 3D multi-marginal OT
9 Conclusions and discussion
We show that the SHAM sort-matching algorithm converges to the solution of the continuous optimal transport problem in the limit of large number of samples. By employing the discrete OT formulation, where the distributions are represented as probability mass functions instead of samples, the matching relation (transport plan) can be found with standard OT solvers at negligible compute and memory cost. We present the SHAM-OT framework, which straightforwardly extends beyond the standard SHAM hard-matching of two distributions: (i) it can naturally include scatter models via regularization, and (ii) it can match more than two distributions in the multi-marginal variant. We show the agreement of OT and SHAM in realistic demonstration examples of matching the halo mass function and the galaxy luminosity function, from both analytical prescription for the halo mass function and from a simulated halo catalogue. We demonstrate how regularized OT can be used to create scatter on the mass-luminosity relation, while satisfying the marginal constraints by construction. Finally, we show that SHAM-OT can be used to match more than two functions, through an example of matching halo mass, galaxy luminosity, and HI mass functions. The SHAM-OT provides instantaneous solutions to the abundance matching problems, which presents multiple benefits for generating simulations, especially if the matching is recomputed frequently for different parameters of luminosity or HI mass functions.
The SHAM-OT framework can be further studied in several directions. The properties of scatter can further be modelled by modifying the metric transport cost. It can potentially be closely linked with existing models of log-normal scatter, as well as be used to straightforwardly explore new models. The regularization strength $\epsilon$ controlling the amount of scatter can be easily included as part of the model and marginalized over a prior in Bayesian inference.
While the entropic penalty is the most commonly used in the OT community, other regularisers can also implemented to create new scatter models. Quadratic OT (Benamou & Brenier, 2000; Nutz, 2024), which penalizes the transport plan by its squared Frobenius norm $\|Q\|_{F}^{2}$, is particularly interesting for this as it has deep connections with fluid dynamics (Benamou & Brenier, 2000), Lagrangian reconstruction methods for large-scale structure (Brenier et al., 2003), and thermal diffusion (Jordan et al., 1998).
Acknowledgements
We acknowledge the use of the following software packages: numpy (Van Der Walt et al., 2011), galsbi (Fischbacher et al., 2024b), POT (Flamary et al., 2021), PyCosmo (Refregier et al., 2018; Tarsitano et al., 2021; Moser et al., 2022; Hitz et al., 2024), scipy (Virtanen et al., 2020). Plots were created using matplotlib (Hunter, 2007) and trianglechain (Fischbacher et al., 2023; Kacprzak & Fluri, 2022).This project was supported in part by grant 200021_192243 from the Swiss National Science Foundation. The authors acknowledge support from SERI and the SKACH consortium.
Software and Data Availability
The code to reproduce this analysis will be made available upon publication under this link: https://cosmo-gitlab.phys.ethz.ch/cosmo_public/sham-ot/. The simulation data can be made available upon request.